An overseas shipment of 5 foreign automobiles contains 2 that have slight paint

blemishes. If an agency receives 3 of these automobiles at random, find the

probability distribution of the random variable X representing the number of

automobiles with paint blemishes purchased by the agency. Find the mean

number of automobiles with paint blemishes. Also, calculate the variation.


1
Expert's answer
2022-04-13T18:04:35-0400

Let's assign a number from 1 to 5 to each automobile and let those ones with paint blemishes have numbers 1 and 2

List of all different samples of 3 automobiles out of 5:

1 2 3

1 2 4

1 2 5

1 3 4

1 3 5

1 4 5

2 3 4

2 3 5

2 4 5

3 4 5

We can see that there are one sample ({3, 4, 5}) with no blemished automobiles, 6 samples with one blemished automobile ({1, 3, 4}, {1, 3, 5}, {1, 4, 5}, {2, 3, 4}, {2, 3, 5}, {2, 4, 5}) and 3 samples with 2 blemished automobiles ({1, 2, 3}, {1, 2, 4}, {1, 2, 5}). Which means that P(X = 0) = 1/10 = 0.1, P(X = 1) = 6/10 = 0.6, P(X=2) = 3/10 = 0.3

Mean:

"\mu=\sum_{i=0}^{2}P(X_i)X_i=\\
=0.1\cdot0+0.6\cdot1+0.3\cdot2=1.2"

Variance:

"\sigma^2=\sum_{i=0}^{2}P(X_i)(X_i-\mu)^2=\\
=0.1\cdot1.2^2+0.6\cdot0.2^2+0.3\cdot0.8^2=0.36"

Standard deviation:

"\sigma=\sqrt{\sigma^2}=0.6"

Coefficient of variation:

"CV=\frac{\sigma}{\mu}=0.5"